# 2]International Research Center for Intelligent Perception and Computation, Xi’an 710071, China

### v1

@@ 3.3.2 Phenomenon 2
Figure 3: (a) The horizontal coordinates are the indexes of the classes, and 1 to 9 indicate the classes that are most similar to the class represented by the vertical coordinates to the least similar, respectively. Each element represents the similarity of the geometry between classes. See Appendix D for detailed class names. (b) Same as (a). (c) The inner product between all eigenvectors of dog and all eigenvectors of cat in CIFAR-10. The sum of the first five diagonal elements of $M1$ is equal to the value of the element in the first column of the first row in (b). (d) The inner product between all eigenvectors of dog and automobile in CIFAR-10. The third row represents the results of the experiments using ResNet-50 and VGG-16 as backbone networks. The axes as well as the meanings of the values are consistent with (a) and (b).
@@
The detailed derivation and proof process of the above equation is shown in the Appendix C. Setting $P$ in $f_{P}(\delta)$ to $64$, when $\delta$ is taken as the first five diagonal elements of $M1$ respectively, the calculation result of $f_{P}(\cdot)$ is shown in Figure 4a. Considering only the first five diagonal elements of $M1$, the probability of the situation shown in Figure 3c occurring is almost 00. Not only that, we observed numerous such phenomena (see Appendix D), thus implying that the phenomena we found could hardly have occurred by chance.
@@ 3.3.3 Phenomenon 3
The phenomenon that features distributions of a similar class has similar geometry only occurs when all features are extracted using the same model. Figure 4b shows that there is a low similarity between the geometry of dog computed by two different ResNet18 trained with random initialization. More examples in Appendix E.
@@ 3.3.4 Phenomenon 4
We conducted further experiments on CIFAR-10 as well as its long-tailed version CIFAR-10-LT. In CIFAR-10-LT, airplane, automobile, bird, and cat are considered head classes and the remaining classes are tail classes. As shown in Figure 4d, we confirmed that the most similar class to the tail class usually belongs to the head class [5] and found that if a tail class and a head class show high similarity in CIFAR-10-LT, they also show high similarity on CIFAR-10.
@@ 4.1 Feature Uncertainty Representation
The sample covariance matrix of class $h$ is estimated as $\Sigma_{h}=\frac{1}{N_{h}}z_{h}z_{h}^{T}\in\mathbb{R}^{P\times P}$. The eigenvalues of the matrix $\Sigma_{h}$ are denoted as $[\lambda_{h}^{1},\dots,\lambda_{h}^{P}]\in\mathbb{R}^{P}$, where $\lambda_{h}^{1}\geq\cdots\geq\lambda_{h}^{P}$. The eigenvector $[\xi_{h}^{1},\dots,\xi_{h}^{P}]\in\mathbb{R}^{P\times P}$, which corresponds one-to-one with the eigenvalues, anchors the geometry of the class $h$ feature distribution, where $\|\xi_{h}^{i}\|_{2}=1,i=1,\dots,P$. Since the distributions of similar class have similar geometry, we propose to represent the uncertainty of $z_{t}^{i}$ by centering a single feature embedding $z_{t}^{i}$ of the tail class $t$ and performing a random translation to $z_{t}^{i}$ along a random linear combination of $\xi_{h}^{1},\dots,\xi_{h}^{P}$. Considering that the “scope" of the distribution is larger in the direction with larger eigenvalues [57], an additional weight $\lambda_{h}^{i}$ is assigned to $\xi_{h}^{i}(i=1,\dots,P)$ when the eigenvectors are randomly combined, which means that $z_{t}^{i}$ is translated farther with higher probability in the direction with larger eigenvalues. In summary, the final form of the proposed method can be represented as
@@ 4.2 Training Scheme
  • Phase-3: Fine-tuned feature sub-network. Fine-tuning the decision boundary can improve the performance of the tail class while compromising the performance of the head class [49]. This is because the feature sub-network is not well adapted to the new decision boundary. Therefore, we propose to freeze the parameters of $g(z,\theta_{2})$ at Phase-3 and fine-tune $f(x,\theta_{1})$ with the original long-tailed data.
@@ 5.1 Datasets and Evaluation Metrics
We evaluate the effectiveness and generalizability of our approach at CIFAR-10/100-LT [2, 15], ImageNet-LT [23], iNaturalist 2018 [37] and OIA-ODIR [18]. For a fair comparison, the training and test images of all datasets are officially split [48, 52, 24], and the Top-1 accuracy on the test set is utilized as a performance metric.
  • Both CIFAR-10 and CIFAR-100 [15] contain $60,000$ images, of which $50,000$ are used for training and $10,000$ for validation, and they contain $10$ and $100$ classes, respectively. For a fair comparison, we use the long-tailed version of the CIFAR dataset. The imbalance factor (IF) is defined as the value of the number of the most frequent class training samples divided by the number of the least frequent class training samples. The imbalance factors we employ in our experiments are $10$, $50$, $100$, and $200$.
  • ImageNet-LT is an artificially produced unbalanced dataset utilizing its balanced version (ImageNet-LT-2012 [33]). It with an imbalance factor of $256$, contains $1000$ classes totaling $115.8k$ images, with a maximum of $1280$ images and a minimum of $5$ images per class.
  • The iNaturalist 2018 dataset is a large-scale real-world dataset that exhibits a long tail. It contains $437,513$ training samples from $8,142$ classes with an imbalance factor of $500$ and three validation samples per class.
  • The OIA-ODIR dataset was made public in $2019$, and it contains a total of $10,000$ fundus images in $8$ classes. See Section 5.6 for a more detailed description.
@@
BBN [56] CVPR 2020 - 79.8 82.1 88.3 - 42.5 47.0 59.1
UniMix [47] NeurIPS 2021 78.5 82.8 84.3 89.7 42.1 45.5 51.1 61.3
MetaSAug [19] CVPR 2021 76.8 80.5 84.0 89.4 39.9 46.8 51.9 61.7
MiSLAS [55] CVPR 2021 - 82.1 85.7 90.0 - 47.0 52.3 63.2
CDB-W-CE [34] IJCV 2022 - - - - - 42.6 - 58.7
GCL [17] CVPR 2022 79.0 82.7 85.5 - 44.9 48.7 53.6 -
RIDE + CR [26] CVPR 2023 - - - - - 50.7 54.3 61.4
OFA [5] ECCV 2020 75.5 82.0 84.4 91.2 41.4 48.5 52.1 65.3
M2m [14] CVPR 2020 - 78.3 - 87.9 - 42.9 - 58.2
RSG [39] CVPR 2021 - 79.6 82.8 - - 44.6 48.5 -
CMO [30] CVPR 2022 - - - - - 50.0 53.0 60.2
FDC [25] TMM 2023 79.7 83.4 86.5 90.6 45.8 50.6 54.1 61.3
@@
Table 3: Top-1 accuracy (%) of ResNext-50 [21] on ImageNet-LT and Top-1 accuracy (%) of ResNet-50 [22] on iNaturalist 2018 for classification. The best and the second-best results are shown in underline bold and bold, respectively. FUR-Decoupled indicates a FUR with decoupled training, and FUR Default indicates a FUR with three-stage training scheme.
@@
BBN [56] CVPR 2020 43.3 45.9 43.7 44.7 49.4 70.8 65.3 66.3
DisAlign [51] CVPR 2021 59.9 49.9 31.8 52.9 68.0 71.3 69.4 70.2
UniMix [47] NeurIPS 2021 - - - 48.4 - - - 69.2
MetaSAug [19] CVPR 2021 - - - 47.3 - - - 68.7
MiSLAS [55] CVPR 2021 65.3 50.6 33.0 53.4 73.2 72.4 70.4 71.6
CDB-W-CE [34] IJCV 2022 - - - 38.5 - - - -
GCL [17] CVPR 2022 - - - 54.9 - - - 72.0
DSB-LADE [24] ICLR 2023 62.6 50.4 33.6 53.2 72.3 70.7 65.8 70.5
LDAM + CR [26] CVPR 2023 60.8 50.3 33.6 52.4 69.3 66.7 61.9 65.7
OFA [5] ECCV 2020 47.3 31.6 14.7 35.2 - - - 65.9
RSG [39] CVPR 2021 63.2 48.2 32.3 51.8 - - - 70.2
GistNet [21] ICCV 2021 52.8 39.8 21.7 42.2 - - - 70.8
BS + CMO [30] CVPR 2022 62.0 49.1 36.7 52.3 68.8 70.0 72.3 70.9
@@ 5.2 Implementation Details
Following the accepted settings [6, 51, 55], the batch sizes on ImageNet-LT and iNaturalist 2018 were taken to be $256$ and $512$, respectively. For a fair comparison, we are consistent with OFA [5] and take $N_{A}$ to be $3$, so $N_{T}$ is $32$ and $64$ on ImageNet-LT and iNaturalist 2018, respectively. More details of the experimental setup are listed in Table 2. The experimental setup on OIA-ODIR will be presented separately in Section 5.6. We trained models on CIFAR-10-LT, CIFAR-100-LT and OIA-ODIR using a single NVIDIA 2080Ti GPU and on ImageNet-LT and iNaturalist 2018 using $4$ NVIDIA 2080Ti GPUs.
@@ 5.3 Comparative Methods
  • Classical and latest long-tailed knowledge transfer methods, include OFA [5], M2m [14], RSG [39], GistNet [21], CMO [30] and FDC [25].
  • Other state-of-the-art methods. They include the two-stage MiSLAS [55], DisAlign [51], BBN [56] with two branches, and the non-transfer augmentation methods UniMix [47], MetaSAug [19], CDB [34], GCL [17] DSB [24] and CR [26].
@@ 5.4 Results on CIFAR-10-LT and CIFAR-100-LT
Compared to CMO, which randomly pastes the image foreground of the tail class onto the background of the head class image, our proposed FUR relies on the observed prior knowledge to recover the underlying distribution of the tail class. GCL constructs the same “feature cloud" for each feature of the tail class to adjust the model logit, without taking into account the differences in domain characteristics between classes. As a result, FUR outperforms similar methods on multiple datasets.
@@ 5.6 Results on OIA-ODIR
The OIA-ODIR dataset contains a total of $10,000$ fundus images in $8$ classes. The eight classes are Normal (N), hypertensive retinopathy (D), glaucoma (G), cataract (C), agerelated macular degeneration (A), hypertension complication (H), pathologic myopia (M), other disease/abnormality (O). Considering that O usually appears together with other diseases, to reduce ambiguity, we adopt the data splitting scheme of [24], using only the data of the first $7$ classes, and the number of training samples and test samples for each class is shown in Figure 5.
@@
We used ResNet-50 as the backbone network. An adam optimizer with a learning rate of $0.1$ (linear decay), a momentum of $0.9$, and a weight decay factor of $0.005$ was adopted to train all networks. In keeping with [24], average precision (AP) was used as the performance metric of the model. We implemented BBN, OFA, and CMO on the OIA-ODIR dataset and compared them with our method. The experimental results (Figure 5) indicate that FUR outperforms the other three methods.

### v2

@@ 3.3.2 Phenomenon 2
Figure 3: (a) The horizontal coordinates are the indexes of the classes, and 1 to 9 indicate the classes that are most similar to the class represented by the vertical coordinates to the least similar, respectively. Each element represents the similarity of the geometry between classes. See Appendix Appendix D for detailed class names. (b) Same as (a). (c) The inner product between all eigenvectors of dog and all eigenvectors of cat in CIFAR-10. The sum of the first five diagonal elements of $M1$ is equal to the value of the element in the first column of the first row in (b). (d) The inner product between all eigenvectors of dog and automobile in CIFAR-10. The third row represents the results of the experiments using ResNet-50 and VGG-16 as backbone networks. The axes as well as the meanings of the values are consistent with (a) and (b).
@@
The detailed derivation and proof process of the above equation is shown in the Appendix Appendix C. Setting $P$ in $f_{P}(\delta)$ to $64$, when $\delta$ is taken as the first five diagonal elements of $M1$ respectively, the calculation result of $f_{P}(\cdot)$ is shown in Figure 4a. Considering only the first five diagonal elements of $M1$, the probability of the situation shown in Figure 3c occurring is almost 00. Not only that, we observed numerous such phenomena (see Appendix Appendix D), thus implying that the phenomena we found could hardly have occurred by chance.
@@ 3.3.3 Phenomenon 3
The phenomenon that features distributions of a similar class has similar geometry only occurs when all features are extracted using the same model. Figure 4b shows that there is a low similarity between the geometry of dog computed by two different ResNet18 trained with random initialization. More examples in Appendix Appendix E.
@@ 3.3.4 Phenomenon 4
We conducted further experiments on CIFAR-10 as well as its long-tailed version CIFAR-10-LT. In CIFAR-10-LT, airplane, automobile, bird, and cat are considered head classes and the remaining classes are tail classes. As shown in Figure 4d, we confirmed that the most similar class to the tail class usually belongs to the head class paper5 and found that if a tail class and a head class show high similarity in CIFAR-10-LT, they also show high similarity on CIFAR-10.
@@ 4.1 Feature Uncertainty Representation
The sample covariance matrix of class $h$ is estimated as $\Sigma_{h}=\frac{1}{N_{h}}z_{h}z_{h}^{T}\in\mathbb{R}^{P\times P}$. The eigenvalues of the matrix $\Sigma_{h}$ are denoted as $[\lambda_{h}^{1},\dots,\lambda_{h}^{P}]\in\mathbb{R}^{P}$, where $\lambda_{h}^{1}\geq\cdots\geq\lambda_{h}^{P}$. The eigenvector $[\xi_{h}^{1},\dots,\xi_{h}^{P}]\in\mathbb{R}^{P\times P}$, which corresponds one-to-one with the eigenvalues, anchors the geometry of the class $h$ feature distribution, where $\|\xi_{h}^{i}\|_{2}=1,i=1,\dots,P$. Since the distributions of similar class have similar geometry, we propose to represent the uncertainty of $z_{t}^{i}$ by centering a single feature embedding $z_{t}^{i}$ of the tail class $t$ and performing a random translation to $z_{t}^{i}$ along a random linear combination of $\xi_{h}^{1},\dots,\xi_{h}^{P}$. Considering that the “scope” of the distribution is larger in the direction with larger eigenvalues paper51 , an additional weight $\lambda_{h}^{i}$ is assigned to $\xi_{h}^{i}(i=1,\dots,P)$ when the eigenvectors are randomly combined, which means that $z_{t}^{i}$ is translated farther with higher probability in the direction with larger eigenvalues. In summary, the final form of the proposed method can be represented as
@@ 4.2 Training Scheme
  • Phase-3: Fine-tuned feature sub-network. Fine-tuning the decision boundary can improve the performance of the tail class while compromising the performance of the head class paper43 . This is because the feature sub-network is not well adapted to the new decision boundary. Therefore, we propose to freeze the parameters of $g(z,\theta_{2})$ at Phase-3 and fine-tune $f(x,\theta_{1})$ with the original long-tailed data.
@@ 5.1 Datasets and Evaluation Metrics
We evaluate the effectiveness and generalizability of our approach at CIFAR-10/100-LT paper2 ; paper15 , ImageNet-LT paper22 , iNaturalist 2018 paper32 and OIA-ODIR benchmark . For a fair comparison, the training and test images of all datasets are officially split paper42 ; paper46 ; paper53 , and the Top-1 accuracy on the test set is utilized as a performance metric.
  • Both CIFAR-10 and CIFAR-100 paper15 contain $60,000$ images, of which $50,000$ are used for training and $10,000$ for validation, and they contain $10$ and $100$ classes, respectively. For a fair comparison, we use the long-tailed version of the CIFAR dataset. The imbalance factor (IF) is defined as the value of the number of the most frequent class training samples divided by the number of the least frequent class training samples. The imbalance factors we employ in our experiments are $10$, $50$, $100$, and $200$.
  • ImageNet-LT is an artificially produced unbalanced dataset utilizing its balanced version (ImageNet-LT-2012 paper28 ). It with an imbalance factor of $256$, contains $1000$ classes totaling $115.8k$ images, with a maximum of $1280$ images and a minimum of $5$ images per class.
  • The iNaturalist 2018 dataset is a large-scale real-world dataset that exhibits a long tail. It contains $437,513$ training samples from $8,142$ classes with an imbalance factor of $500$ and three validation samples per class. The OIA-ODIR dataset was made public in $2019$, and it contains a total of $10,000$ fundus images in $8$ classes. See Section 5.6 for a more detailed description.
@@
BBN paper50 CVPR 2020 - 79.8 82.1 88.3 - 42.5 47.0 59.1
UniMix paper41 NeurIPS 2021 78.5 82.8 84.3 89.7 42.1 45.5 51.1 61.3
MetaSAug paper18 CVPR 2021 76.8 80.5 84.0 89.4 39.9 46.8 51.9 61.7
MiSLAS paper49 CVPR 2021 - 82.1 85.7 90.0 - 47.0 52.3 63.2
CDB-W-CE paper29 IJCV 2022 - - - - - 42.6 - 58.7
GCL paper17 CVPR 2022 79.0 82.7 85.5 - 44.9 48.7 53.6 -
RIDE + CR paper52 CVPR 2023 - - - - - 50.7 54.3 61.4
OFA paper5 ECCV 2020 75.5 82.0 84.4 91.2 41.4 48.5 52.1 65.3
M2m paper14 CVPR 2020 - 78.3 - 87.9 - 42.9 - 58.2
RSG paper34 CVPR 2021 - 79.6 82.8 - - 44.6 48.5 -
CMO paper26 CVPR 2022 - - - - - 50.0 53.0 60.2
FDC paper55 TMM 2023 79.7 83.4 86.5 90.6 45.8 50.6 54.1 61.3
@@
Table 3: Top-1 accuracy (%) of ResNext-50 paper20 on ImageNet-LT and Top-1 accuracy (%) of ResNet-50 paper21 on iNaturalist 2018 for classification. The best and the second-best results are shown in underline bold and bold, respectively. FUR-Decoupled indicates a FUR with decoupled training, and FUR Default indicates a FUR with three-stage training scheme.
@@
BBN paper50 CVPR 2020 43.3 45.9 43.7 44.7 49.4 70.8 65.3 66.3
DisAlign paper45 CVPR 2021 59.9 49.9 31.8 52.9 68.0 71.3 69.4 70.2
UniMix paper41 NeurIPS 2021 - - - 48.4 - - - 69.2
MetaSAug paper18 CVPR 2021 - - - 47.3 - - - 68.7
MiSLAS paper49 CVPR 2021 65.3 50.6 33.0 53.4 73.2 72.4 70.4 71.6
CDB-W-CE paper29 IJCV 2022 - - - 38.5 - - - -
GCL paper17 CVPR 2022 - - - 54.9 - - - 72.0
DSB-LADE paper53 ICLR 2023 62.6 50.4 33.6 53.2 72.3 70.7 65.8 70.5
LDAM + CR paper52 CVPR 2023 60.8 50.3 33.6 52.4 69.3 66.7 61.9 65.7
OFA paper5 ECCV 2020 47.3 31.6 14.7 35.2 - - - 65.9
RSG paper34 CVPR 2021 63.2 48.2 32.3 51.8 - - - 70.2
GistNet paper20 ICCV 2021 52.8 39.8 21.7 42.2 - - - 70.8
BS + CMO paper26 CVPR 2022 62.0 49.1 36.7 52.3 68.8 70.0 72.3 70.9
@@ 5.2 Implementation Details
Following the accepted settings paper6 ; paper45 ; paper49 , the batch sizes on ImageNet-LT and iNaturalist 2018 were taken to be $256$ and $512$, respectively. For a fair comparison, we are consistent with OFA paper5 and take $N_{A}$ to be $3$, so $N_{T}$ is $32$ and $64$ on ImageNet-LT and iNaturalist 2018, respectively. More details of the experimental setup are listed in Table 2. The experimental setup on OIA-ODIR will be presented separately in Section 5.6. We trained models on CIFAR-10-LT, CIFAR-100-LT and OIA-ODIR using a single NVIDIA 2080Ti GPU and on ImageNet-LT and iNaturalist 2018 using $4$ NVIDIA 2080Ti GPUs.
@@ 5.4 Results on CIFAR-10-LT and CIFAR-100-LT
Compared to CMO, which randomly pastes the image foreground of the tail class onto the background of the head class image, our proposed FUR relies on the observed prior knowledge to recover the underlying distribution of the tail class. GCL constructs the same “feature cloud” for each feature of the tail class to adjust the model logit, without taking into account the differences in domain characteristics between classes. As a result, FUR outperforms similar methods on multiple datasets.
@@ 5.6 Results on OIA-ODIR
The OIA-ODIR dataset contains a total of $10,000$ fundus images in $8$ classes. The eight classes are Normal (N), hypertensive retinopathy (D), glaucoma (G), cataract (C), agerelated macular degeneration (A), hypertension complication (H), pathologic myopia (M), other disease/abnormality (O). Considering that O usually appears together with other diseases, to reduce ambiguity, we adopt the data splitting scheme of paper53 , using only the data of the first $7$ classes, and the number of training samples and test samples for each class is shown in Figure 5.
@@
We used ResNet-50 as the backbone network. An adam optimizer with a learning rate of $0.1$ (linear decay), a momentum of $0.9$, and a weight decay factor of $0.005$ was adopted to train all networks. In keeping with paper53 , average precision (AP) was used as the performance metric of the model. We implemented BBN, OFA, and CMO on the OIA-ODIR dataset and compared them with our method. The experimental results (Figure 5) indicate that FUR outperforms the other three methods.